Suppose that the distance of fly balls hit to the outfield (in baseball) is normally distributed with a mean of 250 feet and a standard deviation of 50 feet. We randomly sample 49 fly balls. (a) What is the probability that the 49 balls traveled an average of less than 240 feet? (Round to 3 decimal places) (b) Find the 80th percentile of the distribution of the average of 49 fly balls.
Answer #1

Solution :

Given that ,

mean = \mu = 250

standard deviation = \sigma = 50

n = 49

\mu\bar x = 250

\sigma\bar x =\sigma  / \sqrt n = 50/ \sqrt 49=7.1429

P(\bar x <240 ) = P[(\bar x - \mu \bar x ) / \sigma \bar x < (240 - 250) / 7.1429]

= P(z <-1.40 )

Using z table

=0.0808

=0.081

(B)

Using standard normal table,

P(Z < z) = 80%

= P(Z < z) = 0.80  

= P(Z <0.84 ) = 0.80

z = 0.84

Using z-score formula  

\bar x = z * \sigma \bar x+ \mu \bar x

\bar x = 0.84*7.1429+250

\bar x = 256.000